Identify the slope and y-intercept of the line with the given equation

y=x-4
y=5x+2
2x+y=-6

Answer by mananth(16946)


compare the equations with y=mx+b where m is the slope and b is the y intercept
y=x-4
m=1 the slope and y intercept = -4
y=5x+2
m=5 and y intercept = 2
2x+y=-6
y=-2x+6
m=-2 y intercept = 6